OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (souvik):

let Tim is x years old and his coach is y years old.. so..y=x+10 3 years ago Tim was (x-3) years old and his coach was (y-3) years old.. 3 years ago..(y-3)=2(x-3) now you can find x and y..